OpenStudy (anonymous):

solve proportion 3:(x-2) x:1

OpenStudy (anonymous):

If you put x = 3 I think you will get it ..

OpenStudy (anonymous):

No problem.. Can you cross multiply ???

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

It is like: \[\frac{3}{x - 2} = \frac{x}{1}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You will get: \[x^2 - 2x = 3\] \[x^2 - 2x - 3 = 0 \implies (x-3)(x+1) = 0\] Find \(x\) from here..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you find x from here ??

OpenStudy (anonymous):

Separately put the brackets to 0..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

It will be better if you will tell what did you get instead of just ok..

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i got x=3 x=-1

OpenStudy (anonymous):

Yes well done you are right..

OpenStudy (anonymous):

thanks alot waterineyes

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

:)